[Hallc_running] [New Logentry] RC Weekly summary and update

munoz at jlab.org munoz at jlab.org
Mon Dec 10 11:35:02 EST 2018

Logentry Text:
<p>Since Hall C meeting is cancelled today, and there might be no beam before the end of my term, I include below an ABU summary of the last week. In general we've been running smoothly with 80.5 h ABU since Wed 3PM, which gives us around 70% efficiency. Only minor issues with SHMS magnets (one IOC and two power supplies power cycles needed over the weekend). A slight DAQ issue (missing reference times) is being investigated today by Eric and Howard.</p><p>The Hall is now in restricted access. The plan is to resume beam by the end of day Wednesday. At that time, we'll proceed with the 4.933 GeV beam runplan (https://redmine.jlab.org/attachments/download/590/runplan_4p9gev.pdf)</p><p>There will be NO RC meeting today. We may meet tomorrow depending on the progress with the beam energy change.</p><p> </p><table border="0"><caption> Week ABU summary</caption><tbody><tr><td>Shift</td><td>ABU (h)</td><td>Comment</td></tr><tr><td><strong>Wed Day</strong></td><td>7.3</td><td>&nbsp
 ;</td></tr><tr><td style="text-align: right;">Swing</td><td>6.5</td><td> </td></tr><tr><td style="text-align: left;"><strong>Thu Owl</strong></td><td>7.1</td><td> </td></tr><tr><td style="text-align: right;">Day</td><td>7.4</td><td> </td></tr><tr><td style="text-align: right;"> Swing</td><td>2.7</td><td>Acc vacuum problems </td></tr><tr><td> <strong>Frid Owl</strong></td><td>6.6</td><td> </td></tr><tr><td style="text-align: right;">Day </td><td>4.1</td><td>Acc optics work for B/D</td></tr><tr><td style="text-align: right;">Swing </td><td>6.1</td><td> </td></tr><tr><td style="text-align: left;"><strong>Sat Owl </strong></td><td>3.5</td><td>Acc troubleshooting beam to B/D</td></tr><tr><td style="text-align: right;">Day </td><td>5.7</td><td>SHMS Q3 IOC reboot (access)</td></tr><tr><td style="text-align: right;">Swing </td><td>7.5</td><td> </td></tr><tr><td style="text-align: left;"><strong>Sun Owl </strong></td>
 <td>5.8</td><td> </td></tr><tr><td styl!
 e="text-align: right;">Day </td><td>7.2</td><td> </td></tr><tr><td style="text-align: right;">Swing </td><td>4.5</td><td>Acc+Hall power supply issues (storm) </td></tr><tr><td style="text-align: left;"><strong>Mon Owl </strong></td><td>5.8/6</td><td>Beam stopped at 6AM</td></tr></tbody></table>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3637037
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181210/14d8242a/attachment-0001.html>

More information about the Hallc_running mailing list